OpenStudy (meganhall0517):

. Amber owns a business. She charges $24.95 for each CD she makes.. Her expenses include $575 for equipment, and $0.80 for each blank CD. Which of these equations could be used to calculate her profit p for n CD’s? a. p = (24.95 – 0.8)n – 575 b. p = (24.95 + 0.8)n + 575 c. p = 24.95n – 574.2 d. p = 24.95n + 575

OpenStudy (studygurl14):

She charges 24.95 per CD, and it costs her 0.8 per cd, so you have (24.95-0.8)n, where n= amount of cds then you'd subtract 575 because it is an expense p = profit So (24.95-0.8)n - 575 = p

OpenStudy (meganhall0517):

i have no clue how to do this

OpenStudy (meganhall0517):

so a

OpenStudy (sissyedgar):

I agree with Sturygirl14!! ♥

OpenStudy (meganhall0517):

Which equation has a solution of -5? a. 2a – 6 = 4 b. 3a + 7 = 8 c. 3a – 7 = 2 d. 3a + 19 = 16 4 5

OpenStudy (meganhall0517):

can someone help

OpenStudy (sissyedgar):

It would be better to ask in a new tread. x) I can still help you tho. ♥